From 1.0 M NH4OH and 1.0 M NH4Cl reagent stock solutions prepare 100 mL of the solutions described in the table and complete the table.

Volume NH3 Volume NH4+ Solução [NH3]/[NH4+][NH3] + [NH4+] 1:1 0,10 1:1 0,20 1:10 0,20 10:1 0,20
Answer #1

1.

[NH3] /[NH4+] = 1:1

[NH3]= [NH4+]

And [NH3] + [NH4+] = 10

Therefore

[NH3]+ [NH3]= 10

[NH3]= 5

And [NH4+]=5 ml

2.

[NH3] /[NH4+] = 1:1

[NH3]= [NH4+]

And [NH3] + [NH4+] = 20

Therefore

[NH3]+ [NH3]= 20

[NH3]= 10 ml

And [NH4+]=10 ml

3.

[NH3] /[NH4+] = 1:10

10[NH3]= [NH4+]

And [NH3] + [NH4+] = 20

Therefore

10 [NH3] + [NH3] = 20

11 [NH3]=20

[NH3]=1.82 ml

And

[NH3]+ [NH4+]= 20

[NH4+]= 20-1.82 ml

= 18.2

Cross check:

[NH3] /[NH4+] = 1:10

1.82/18.2=1:10

4.

[NH3] /[NH4+] = 10:1

[NH3]=10 [NH4+]

And [NH3] + [NH4+] = 20

Therefore

10 [NH4+] + [NH4+] = 20

11 [NH4+]=20

[NH4+]=1.82 ml

And

[NH3]+ [NH4+]= 20

[NH3]= 20-1.82 ml

= 18.2

Cross check:

[NH3] /[NH4+] = 1:10

18.2/1.82=1:10

1.

[NH3] /[NH4+] = 1:1

[NH3]= [NH4+]

And [NH3] + [NH4+] = 10

Therefore

[NH3]+ [NH3]= 10

[NH3]= 5

And [NH4+]=5 ml

2.

[NH3] /[NH4+] = 1:1

[NH3]= [NH4+]

And [NH3] + [NH4+] = 20

Therefore

[NH3]+ [NH3]= 20

[NH3]= 10 ml

And [NH4+]=10 ml

3.

[NH3] /[NH4+] = 1:10

10[NH3]= [NH4+]

And [NH3] + [NH4+] = 20

Therefore

10 [NH3] + [NH3] = 20

11 [NH3]=20

[NH3]=1.82 ml

And

[NH3]+ [NH4+]= 20

[NH4+]= 20-1.82 ml

= 18.2

Cross check:

[NH3] /[NH4+] = 1:10

1.82/18.2=1:10

4.

[NH3] /[NH4+] = 10:1

[NH3]=10 [NH4+]

And [NH3] + [NH4+] = 20

Therefore

10 [NH4+] + [NH4+] = 20

11 [NH4+]=20

[NH4+]=1.82 ml

And

[NH3]+ [NH4+]= 20

[NH3]= 20-1.82 ml

= 18.2

Cross check:

[NH3] /[NH4+] = 1:10

18.2/1.82=1:10